# On the role of overparametrization in Quantum Approximate Optimization

### v1

@@ Abstract
Variational quantum algorithms have emerged as a cornerstone of contemporary quantum algorithms research. While they have demonstrated considerable promise in solving problems of practical interest, efficiently determining the minimal quantum resources necessary to obtain such a solution remains an open question. In this work, inspired by concepts from classical machine learning, we investigate the impact of overparameterization on the performance of variational algorithms. Our study focuses on the quantum approximate optimization algorithm (QAOA)—a prominent variational quantum algorithm designed to solve combinatorial optimization problems. We investigate if circuit overparametrization is necessary and sufficient to solve such problems in QAOA, considering two representative problems — MAX-CUT and MAX-2-SAT. For MAX-CUT we observe that overparametriation is both sufficient and (statistically) necessary for attaining exact solutions, as confirmed numerically for up to $20$ qubits. In fact, for MAX-CUT on 2-regular graphs we show the necessity to be exact, based on the analytically found optimal depth. In sharp contrast, for MAX-2-SAT, underparametrized circuits suffice to solve most instances. This result highlights the potential of QAOA in the underparametrized regime, supporting its utility for current noisy devices.
@@ I Introduction
While variational quantum algorithms (VQAs) have shown immense promise in addressing problems of practical significance [11, 12, 13, 14, 15, 16], determining the minimal resources required to solve a specific problem remains a challenge. In the context of VQAs these resources are typically quantified by the number of required parameterized gates. A similar question has been explored in the context of classical neural networks (CNNs), where the resources are typically quantified by the number of model parameters. It has been shown that overparametrized models — models with more parameters than needed to represent the training data — may partially address this issue by finding solutions that can reduce both training and generalization errors [17, 18]. Motivated by the similarities that the VQAs and CNNs share, the impact of overparametrization on VQAs has been investigated, with promising results. It has been shown to improve the success probability and convergence rate [19] of VQE, help in avoiding spin-glass-like cost landscapes [20, 21] and assist in mitigating the effect of noise [22]. More strikingly, in connection to the question of required quantum resources, overparametrization was shown to improve the probability of finding solutions to certain problems with specific problem dependent ansatz [19].
@@
Despite showing promise in improving VQA performance, reaching the overparametrization regime can often require circuit depth that exceed the capabilities of NISQ devices [23, 19]. Moreover, such large circuit depths may hinder trainability by inducing barren plateaus in the cost landscape. Motivated by this, we study if overparametrization is (i) sufficient and (ii) necessary for attaining solutions with a high probability for a given problem and a variational ansatz. We accomplish this by investigating the relationship between the circuit overparametrization depth and the optimal depth — depths required to achieve overparametrization and to find a solution to a given problem, respectively. In this work, we address this question with a focus on a special class of VQAs — Quantum Approximate Optimization Algorithm (QAOA). It is an algorithm designed to solve combinatorial optimization problems on near-term quantum processors, utilizing a problem dependent ansatz. With long enough circuits, QAOA is known to approximate the adiabatic evolution, thus guaranteeing success [24, 25, 26]. However no such guarantee is known for intermediate depth QAOA sequences. Thus, establishing a relationship between the optimal QAOA depth and the overparametrization depth, may offer valuable insights into performance guarantees for QAOA in the intermediate depth regime.
@@
In this work we benchmark the performance of a QAOA sequence with respect to MAX-CUT problem on regular and random unweighted graphs as well as with random instances of the MAX-2-SAT problem. For the MAX-CUT problem on 2-regular graphs (a.k.a. rings of disagrees), we analytically establish that optimal QAOA depth scales as $\lfloor\frac{n}{2}\rfloor$ for even number of qubits $n$. For odd $n$ the same scaling is confirmed numerically. Importantly, we observe that overparametrization depth exactly coincides with the optimal circuit depth for up to 20 qubits. We numerically demonstrate that this agreement approximately holds for the MAX-CUT problem on unweighted regular graphs as well as unweighted random graphs: the solutions are likely to be found exactly at (or in a close vicinity of) the overparametrization depth. The case of MAX-2-SAT problem, however, demonstrates a different behavior. We find that the solutions to MAX-2-SAT instances could be obtained at depths much smaller than the overparametrized depth, with a high probability. These observations demonstrate that while overparametrization is sufficient to obtain solutions with a high probability, it may not be a necessary condition across all problems considered; while it is statistically necessary for MAX-CUT, it is not so in the case for MAX-2-SAT.
@@ II.1 Quantum Approximate Optimization Algorithm
The challenge in determining the optimal depth is further highlighted by the critical role that the initial choice of $\bm{\beta}$ and $\bm{\gamma}$ plays in the optimization process. Convergence to the desired accuracy heavily depends on these initial parameters. The ease of finding a solution can be quantified by the success probability, defined as a fraction of successful optimization runs (see Appendix D for details), where each run starts from different initial choices of $\bm{\beta}$ and $\bm{\gamma}$.
@@ II.2 Overparametrization in variational circuits
In this section, we investigate this question in the context of QAOA, focusing on instances of the MAX-CUT and the MAX-2-SAT problems. The details of these problems and their Hamiltonian formulations are provided in Appendix A. In particular, we investigate if overparametrization is a sufficient condition for a QAOA sequence to achieve optimal solutions and, conversely, whether it is necessary for attaining such solutions. These questions are particularly significant since both $p_{c}$ and $p^{*}$ are inherently dependent on the problem Hamiltonian, owing to the problem-dependent ansatz used in QAOA.
@@ III.1 MAX-CUT on $2$-regular graphs
Interestingly, previous numerical studies on the ring of disagrees have reported that the optimal depth also scales as $p^{*}=\lfloor\frac{n}{2}\rfloor$ [44], which has been conjectured to hold for any system size $n$. Indeed, we numerically reproduce this result for up to $n=20$ qubits, reaching the tolerance $\epsilon\sim 10^{-12}$. This is demonstrated in figure 1 for $n=20$ (blue circles), where a sharp drop in $E^{*}_{p}(H)-E_{g}$ is observed at $p=10$. The observed drop in $E^{*}_{p}(H)-E_{g}$ is characteristic to the problem and in Appendix B we argue that the drop remains finite even in the thermodynamic limit for even values of $n$. For odd values of $n$ however the drop can be shown to vanish in the limit $n\rightarrow\infty$. Building on our numerical observations, we provide an analytical guarantee for the optimal QAOA circuit depth.
@@ III.2 MAX-CUT on random graphs
To further substantiate the role of overparametrization, we systematically investigate the performance of QAOA across a broader class NP-hard MAX-CUT problem for graph sizes of up to $n=10$ nodes. We consider both $k$-regular graphs—a natural generalization of the 2-regular graphs considered in the previous section— with $k\in[3,7]$ and unweighted random graphs. For both cases, qualitatively similar results are obtained. Thus, the section focuses on the results obtained for the unweighted random graphs, which represent a richer class of problems.
@@
For each graph size $n$ and each edge probability $q\in[0.3,0.9]$ varied with a step size of $0.1$, we sample $50$ random instances. Each instance is then minimized using QAOA with the layerwise heuristic strategy (see Appendix D) using $20$ optimization runs with random parameter initializations. Figure 2 illustrates different types of energy behavior with respect to circuit depth. The blue curve depicts the case when the solution is found exactly at the overparametrization depth $p_{c}$. The green and red lines, on the other hand, illustrate that overparametrization is not necessary for preparing the exact ground state for certain problem instances. However, a systematic comparison of the optimal and overparametrization circuit depth demonstrates that most of the instances are solved exactly at or slightly before the overparametrization depth $p_{c}$ (similar to the blue curve in figure 2). Indeed, this is demonstrated in the blue curve of figure 3(a) which indicates an abrupt increase in the fraction of solved instances, as one transitions from the underparametrized to the overparameterized regime. Note that, here the normalized depth $p/p_{c}$ is used as it enables a uniform comparison of QAOA performance across problem instances with varying overparameterization depths.
@@
The importance of overparametrization becomes clearer if success probability is taken into account. To accomplish this, for each value of $p/p_{c}$ we compute the average success probability over all instances that were solved by the considered normalized depth. The blue curve in figure  3(b) demonstrates the variation of this success probability with respect to the normalized circuit depth. One can clearly observe that the success probability experiences an abrupt increase when transitioning from the underparametrized to the overparametrized regime. Combining the observations from figures 3(a) and (b) we conclude that overparametrization is sufficient to solve a MAX-CUT instance with a high probability. Conversely, while overparametrization is not a necessary condition, a major proportion of the problem instances get solved solely in the immediate vicinity of $p/p_{c}=1$. Moreover, the success probability of finding the solutions experiences a significant improvement only as one transitions into the overparametrized regime.
@@
Figure 3: Optimization results for MAX-CUT (in blue) and MAX-2-SAT (in red) on $n=7$ qubit instances. (a) Fraction of solved instances with respect to the normalized circuit depth $p/p_{c}$. The fraction is computed over all the considered instances (50 random instances for each edge probability $q$, and 50 random instances for each number of clauses $m$, for MAX-CUT and MAX-2-SAT, respectively). (b) Fraction of converged optimization runs for the instances considered in (a) with respect to the normalized circuit depth.
@@ III.3 MAX-2-SAT
In this section we investigate how the observed relationship between the overparametrization and optimal circuit depth generalizes to other combinatorial problems. To this end we turn our attention to a different class of NP-hard combinatorial optimization problems — MAX-2-SAT. A detailed account on the MAX-2-SAT problem and its embedding into a QAOA problem Hamiltonian can be found in Appendix A. In the numerical experiments, we consider MAX-2-SAT instances with $n\in[4,10]$ variables and number of clauses $m\in[2,2n]$ with a step size of $2$, for each value of $n$. The choice of $m$ was dictated by the observation that overparametrization depth $p_{c}$ stagnates at the maximum of $2^{n}-1$ for the clause density $\alpha=m/n\gtrsim 2$, with a similar stagnation reported for the optimal circuit depth [45]. $50$ instances of random, uniformly generated MAX-2-SAT problem were considered for each combination of $n$ and $m$. Each instance is minimized using QAOA with the layerwise heuristic strategy using $50$ optimization runs with random parameter initializations.
@@
In a striking contrast to the observation for MAX-CUT, most of the MAX-2-SAT instances were solved well within the underparametrized region, $p/p_{c}\ll 1$. This is demonstrated in the red curve of figure 3(a) which shows the fraction of solved MAX-2-SAT instances with respect to the normalized depth. This result indicates that overparametriazion is not necessary to solve instances of the MAX-2-SAT problem. Importantly, as highlighted by the red curve in figure 3(b) the success probability also becomes significant in the underparametrized regime, even at $p/p_{c}\ll 1$. This is in a stark contrast with our observation for MAX-CUT, where the success probability remained low at $p/p_{c}\ll 1$, even though overparametrization was not deemed necessary to solve the instances.
@@
The above observations suggest that the overparametrization depth serves as a reasonable estimate for the depths required to solve instances of the MAX-CUT problem with a high probability. In contrast, for MAX-2-SAT, the relation between the optimal and the overparametrization depth is less obvious. Motivated by this, we investigate the scaling of both $p^{*}$ and $p_{c}$ with respect to the size of the problem instances $n$. Out of $350$ instances considered in our experiment, only $21$ were not solved by the depth $p=150$; for these instances we set $p^{*}=2^{n}-1$. The results can be seen in figure 4 for problem instances of density $\alpha=m/n=2$. First, it can be observed, that the overparametrization depth $p_{c}$ scales as $O(2^{n})$, thereby saturating the dimension of the Hilbert space. The optimal circuit depth, despite seemingly exhibiting an exponential behavior as well, demonstrates a more modest growth. As a result, the difference between the two is clearly seen to increase exponentially with problem size $n$, thereby confirming the observation that a majority of MAX-2-SAT instances can be solved well within the undeparametrized regime.
@@
IV Conclusion
In classical machine learning, overparameterization has been extensively studied and is known to improve both optimization and generalization performance [17, 18]. These insights have motivated analogous investigations for VQAs, owing to their structural similarities with classical machine learning models such as neural networks. However, despite these studies, the specific effect of overparametrization on a circuit’s ability to solve concrete problems has remained largely unexplored. In this work, we addressed this question in the setting of QAOA with respect to distinct classes of NP-hard combinatorial optimization problems — MAX-CUT and MAX-2-SAT.

### v2

@@ Abstract
Variational quantum algorithms have emerged as a cornerstone of contemporary quantum algorithms research. While they have demonstrated considerable promise in solving problems of practical interest, efficiently determining the minimal quantum resources necessary to obtain such a solution remains an open question. In this work, inspired by concepts from classical machine learning, we investigate the impact of overparametrization on the performance of variational algorithms. Our study focuses on the quantum approximate optimization algorithm (QAOA)—a prominent variational quantum algorithm designed to solve combinatorial optimization problems. We investigate if circuit overparametrization is necessary and sufficient to solve such problems in QAOA, considering two representative problems — MAX-CUT and MAX-2-SAT. For MAX-CUT on 2-regular graphs we observe that overparametrization is both sufficient and necessary. To establish this, we analytically show that the optimal circuit depth for such problems scale as $n/2$ for even $n$. For a more general case of random graphs, the overparametrization was observed to be sufficient, yet necessary only for a statistically dominant fraction of instances. In sharp contrast, for MAX-2-SAT, underparametrized circuits suffice to solve most instances. This result highlights the potential of QAOA in the underparametrized regime, supporting its utility for current noisy devices.
@@ I Introduction
While variational quantum algorithms (VQAs) have shown immense promise in addressing problems of practical significance [25, 46, 54, 14, 1, 2], estimating how the resources scale with respect to the ease of finding a solution remains an open challenge. In this regard, overparametrized quantum circuits, which use a number of parameters greater than a given threshold, become promising by allowing for establishing certain convergence guaranties. Indeed, it has been shown analytically that for certain types of overparametrized ansatze [67] VQE converges to a small error exponentially with the number of gradient descent iterations. Moreover, in this regime, despite exponential suppression in the circuit parameter updates, the gradient descent dynamics still converges efficiently to the solution [40, 7], avoiding the problem of barren plateaus [37]. Overparametrization has also been shown to improve the success probability and convergence rate [36] of VQE, to help in avoiding spin-glass-like cost landscapes [16, 50] and to assist in mitigating the effect of noise [27]. Importantly, overparametrization was shown to improve the probability of finding solutions to certain problems with a problem dependent ansatz [36, 67].
@@
Despite showing promise in improving VQA performance, the overparametrization regime often requires circuit depth that exceed the capabilities of NISQ devices [35, 36]. Moreover, although in certain settings some studies suggest otherwise [34, 40], such large circuit depths may hinder trainability by inducing barren plateaus. Motivated by this, we study if overparametrization is (i) sufficient and (ii) necessary for attaining solutions with a high probability for a given problem and a variational ansatz. It is worth noting here, that the notion of overparametrization varies across the literature; under certain definitions existing analytical results do imply sufficiency [67, 40, 7] although the aspect of necessity remains underexplored. In this work however we follow a different definition of overparametrization [29, 36] using the concept of effective quantum dimension (EQD). We investigate the relationship between the circuit overparametrization depth and the optimal depth — depths required to achieve overparametrization (saturate EQD) and to find a solution to a given problem, respectively. We address this question with a focus on a special class of VQAs — Quantum Approximate Optimization Algorithm (QAOA) [69, 44, 41, 5, 4, 17, 48, 21, 24, 26]. It is an algorithm designed to solve combinatorial optimization problems on near-term quantum processors, utilizing a problem dependent ansatz. With long enough circuits, QAOA is known to approximate the adiabatic evolution, thus guaranteeing success [23, 31, 11]. However no such guarantee is known for intermediate depth QAOA sequences. Thus, establishing a relationship between the optimal QAOA depth and the overparametrization depth, may offer valuable insights into performance guarantees for QAOA in the intermediate depth regime.
@@
We benchmark the performance of a QAOA sequence with respect to MAX-CUT problem on regular and random unweighted graphs as well as with random instances of the MAX-2-SAT problem. For the MAX-CUT problem on 2-regular graphs (a.k.a. rings of disagrees), we analytically establish that optimal QAOA depth scales as $\lfloor\frac{n}{2}\rfloor$ for even number of qubits $n$. For odd $n$ the same scaling is confirmed numerically. Importantly, we observe that overparametrization depth exactly coincides with the optimal circuit depth for up to 20 qubits. We numerically demonstrate that this agreement approximately holds for the MAX-CUT problem on unweighted regular graphs as well as unweighted random graphs: the solutions are likely to be found exactly at (or in a close vicinity of) the overparametrization depth. The case of MAX-2-SAT problem, however, demonstrates a different behavior: we find that the solutions to MAX-2-SAT instances could be obtained at depths much smaller than the overparametrized depth, with a high probability. These results demonstrate that overparametrization is sufficient to obtain solutions with a high probability. However, in sharp contrast to previous observations (e.g. [36, 67]), it is not a necessary condition across all problems considered; while it is statistically necessary for MAX-CUT, it is not so in the case for MAX-2-SAT, thereby highlighting the potential of QAOA in the underparametrized regime.
@@ II.1 Quantum Approximate Optimization Algorithm
Determining the optimal depth is further complicated by the sensitivity to the initial values of $\bm{\beta}$ and $\bm{\gamma}$. Convergence to the desired accuracy heavily depends on these initial parameters. The ease of finding a solution can be quantified by the success probability, defined as a fraction of successful optimization runs (see Appendix C for details), where each run starts from different initial choices of $\bm{\beta}$ and $\bm{\gamma}$.
@@ II.2 Overparametrization in variational circuits
In this section, we investigate this question in the context of QAOA, focusing on instances of the MAX-CUT and the MAX-2-SAT problems. In particular, we investigate if overparametrization is a sufficient condition for a QAOA sequence to achieve optimal solutions and, conversely, whether it is necessary for attaining such solutions. These questions are particularly significant since both $p_{c}$ and $p^{*}$ are inherently dependent on the problem Hamiltonian, owing to the problem-dependent ansatz used in QAOA.
@@ III.1 MAX-CUT on $2$-regular graphs
Interestingly, previous numerical studies on the ring of disagrees have reported that the optimal depth also scales as $p^{*}=\lfloor\frac{n}{2}\rfloor$ [60], which has been conjectured to hold for any system size $n$. Indeed, we numerically reproduce this result for up to $n=20$ qubits, reaching the tolerance $\epsilon\sim 10^{-12}$. This is demonstrated in figure 1 for $n=20$ (blue circles), where a sharp drop in $E^{*}_{p}(H)-E_{g}$ is observed at $p=10$. The observed drop in $E^{*}_{p}(H)-E_{g}$ is characteristic to the problem and in Appendix A we argue that the drop remains finite even in the thermodynamic limit for even values of $n$. For odd values of $n$ however the drop can be shown to vanish in the limit $n\rightarrow\infty$. Building on our numerical observations, we provide an analytical guarantee for the optimal QAOA circuit depth.
@@ III.2 MAX-CUT on random graphs
To further substantiate the role of overparametrization, we systematically investigate the performance of QAOA across a broader class NP-hard MAX-CUT problem for graph sizes of up to $n=10$ nodes. Given an arbitrary graph $G=(V,E)$, the MAX-CUT problem can be encoded into an $n=|V|$ qubit problem Hamiltonian
$H=\sum_{(i,j)\in E}Z_{i}Z_{j},$ (9)
@@
whose ground state encodes the solution to the problem. We consider both $k$-regular graphs—a natural generalization of the 2-regular graphs considered in the previous section— with $k\in[3,7]$ and unweighted random graphs. For both cases of random and regular graphs, qualitatively similar results are obtained. Thus, the section focuses on the results obtained for the unweighted random graphs, which represent a richer class of problems.
@@
For each graph size $n$ and each edge probability $q\in[0.3,0.9]$ varied with a step size of $0.1$, we sample $50$ random instances. Each instance is then minimized using QAOA with the layerwise heuristic strategy (this choice is motivated in Appendix C) using $20$ optimization runs with random parameter initializations. Figure 2 illustrates different types of energy behavior with respect to circuit depth. The blue curve depicts the case when the solution is found exactly at the overparametrization depth $p_{c}$. The green and red lines, on the other hand, illustrate that overparametrization is not necessary for preparing the exact ground state for certain problem instances. However, a systematic comparison of the optimal and overparametrization circuit depth demonstrates that most of the instances are solved exactly at or slightly before the overparametrization depth $p_{c}$ (similar to the blue curve in figure 2). Indeed, this is demonstrated in the solid blue curve of figure 3(a) which indicates an abrupt increase in the fraction of solved instances, as one transitions from the underparametrized to the overparametrized regime. Note that, here the normalized depth $p/p_{c}$ is used as it enables a uniform comparison of QAOA performance across problem instances with varying overparametrization depths.
@@
The importance of overparametrization becomes clearer if success probability is taken into account. To accomplish this, for each value of $p/p_{c}$ we compute the average success probability over all instances that were solved by the considered normalized depth. The solid blue curve in figure  3(b) demonstrates the variation of this success probability with respect to the normalized circuit depth. One can clearly observe that the success probability experiences an abrupt increase when transitioning from the underparametrized to the overparametrized regime. Combining the observations from figures 3(a) and (b) we conclude that overparametrization is sufficient to solve a MAX-CUT instance with a high probability. Conversely, while overparametrization is not a necessary condition, a major portion of the problem instances get solved solely in the immediate vicinity of $p/p_{c}=1$. Moreover, the success probability of finding the solutions experiences a significant improvement only as one transitions into the overparametrized regime.
@@
Nevertheless, it is worth noting that the critical role of overparametrization becomes less pronounced for larger tollerance $\epsilon$ in definition 1. In dashed blue lines in figure 3 we demonstrate how fast the instances and optimization runs converge to the error tolerance $\epsilon=0.1$, which show a dramatic improvement compared to lower tollerance. Note that the solution can now be extracted only probabilistically, as single shot measurement will extract the exact solution with at least $1-\epsilon/\Delta\geq 0.9$ probability. However, this would typically be considered a tolerable cost, as it allows to significantly reduce circuit depth—the main obstacle for quantum computing in the NISQ era
Figure 3: Optimization results for MAX-CUT (in blue) and MAX-2-SAT (in red) on $n=7$ qubit instances. Solid and dashed lines depict the convergence for $\epsilon=10^{-8}$ and $\epsilon=10^{-1}$ energy error thresholds, respectively. (a) Fraction of solved instances with respect to the normalized circuit depth $p/p_{c}$. The fraction is computed over all the considered instances (50 random instances for each edge probability $q$, and 50 random instances for each number of clauses $m$, for MAX-CUT and MAX-2-SAT, respectively). (b) Fraction of converged optimization runs for the instances considered in (a) with respect to the normalized circuit depth.
@@ III.3 MAX-2-SAT
In this section we investigate how the observed relationship between the overparametrization and optimal circuit depth generalizes to other combinatorial problems. To this end we turn our attention to a different class of NP-hard combinatorial optimization problems — MAX-2-SAT. Given a boolean formula written in conjunctive normal form with exactly $k$ literals per clause, a MAX-k-SAT problem consists of fining a variable assignment that maximizes the number of satisfied clauses. The problem can be embedded into a Hamiltonian minimization problem for
$H_{\text{SAT}}=\sum_{j}P(j),$ (10)
where $j$ indexes clauses of an instance, and $P(j)$ is the tensor product of projectors that penalizes bit string assignments that do not satisfy the $j$-th clause. When rewritten in the Pauli basis, (10) becomes a $k$-local Ising type Hamiltonian. In this work we focused on the $k=2$ case.
@@
In the numerical experiments, we consider MAX-2-SAT instances with $n\in[4,10]$ variables and number of clauses $m\in[2,2n]$ with a step size of $2$, for each value of $n$. The choice of $m$ was dictated by the observation that overparametrization depth $p_{c}$ stagnates at the maximum of $2^{n}-1$ for the clause density $\alpha=m/n\gtrsim 2$, with a similar stagnation reported for the optimal circuit depth [4]. $50$ instances of random, uniformly generated MAX-2-SAT problem were considered for each combination of $n$ and $m$. Each instance is minimized using QAOA with the layerwise heuristic strategy using $50$ optimization runs with random parameter initializations.
@@
In a striking contrast to the observation for MAX-CUT, most of the MAX-2-SAT instances were solved well within the underparametrized region, $p/p_{c}\ll 1$. This is demonstrated in the solid red curve of figure 3(a) which shows the fraction of solved MAX-2-SAT instances with respect to the normalized depth. This result indicates that overparametrization is not necessary to solve instances of the MAX-2-SAT problem. Importantly, as highlighted by the solid red curve in figure 3(b) the success probability also becomes significant in the underparametrized regime, even at $p/p_{c}\ll 1$. This is in a stark contrast with our observation for MAX-CUT, where for $\epsilon=10^{-8}$ the success probability remained low at $p/p_{c}\ll 1$, even though overparametrization was not deemed necessary to solve the instances. Expectedly, lowering the threshold $\epsilon$ for MAX-2-SAT improves the convergence even further, see dashed red lines in figure 3.
@@
The above observations suggest that the overparametrization depth serves as a reasonable estimate for the depths required to solve instances of the MAX-CUT problem to a negligible energy error. In contrast, for MAX-2-SAT, the correlation between the optimal and the overparametrization depth is less pronounced. Motivated by this, we investigate the scaling of both $p^{*}$ and $p_{c}$ with respect to the size of the problem instances $n$. Out of $350$ instances considered in our experiment, only $21$ were not solved by the depth $p=150$; for these instances we set $p^{*}=2^{n}-1$. The results can be seen in figure 4 for problem instances of density $\alpha=m/n=2$. First, it can be observed, that the overparametrization threshold $p_{c}$ scales with the dimension of the Hilbert space, as it has been established for other notions of overparametrization [40, 67]. Interestingly, in this case $p_{c}$ almost saturates the dimension, as $p_{c}\sim 2^{n}$. The optimal circuit depth, despite seemingly exhibiting an exponential behavior as well, demonstrates a more modest growth, $p^{*}\sim\sqrt{2^{n}}$ for the energy error threshold $\epsilon=10^{-8}$. As a result, the difference between the two is clearly seen to increase exponentially with problem size $n$, thereby confirming the observation that a majority of MAX-2-SAT instances can be solved well within the underparametrized regime. This observation becomes even more pronounced for a higher energy error threshold, $\epsilon=10^{-1}$, as depicted in red in figure 3. Interestingly, the observed scaling implies that upon problem size increase, the instances are getting solved by increasingly more underparametrized circuits, as $p^{*}/p_{c}\sim 1/\sqrt{2^{n}}$.
@@
IV Conclusion and discussion
In classical machine learning, overparametrization has been extensively studied and is known to improve both approximation and generalization performance [8, 6]. These insights have motivated analogous investigations for VQAs, owing to their structural similarities with classical machine learning models such as neural networks. Existing studies tend to focus on the overparametrized quantum circuits [67, 40, 7] which allow establishing theoretical convergence guaranties, proving sufficiency under certain definitions of overparametrization for solving problems with specific ansatz types. A more quantifiable approach for defining overparametrization was introduced by [36] using the concept of effective quantum dimension [29]. In this work the authors demonstrate empirical evidence for sufficiency as well as necessity of overparametrization, by showing convergence improvement upon transitioning into the overparametrized regime. In our work we adopted the same EQD based definition of overparametrization threshold and investigate if it is indeed both sufficient and necessary for finding exact solution with high probability. We addressed this question in the setting of QAOA with respect to distinct classes of NP-hard combinatorial optimization problems — MAX-CUT and MAX-2-SAT.